OpenStudy (anonymous):

for the sequence an=3n-5,what is the value of a10

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

im taking an exam right now i need help :"(

OpenStudy (solomonzelman):

plug in 10 for n a n =3n-5 a 10 =3(10)-5 a 10 =30-5 a 10 =25 good ?

OpenStudy (anonymous):

thanks